⎛ 2 0 1 ⎞
9.Given that A = , ⎜ k 2 3 ⎟ what is the value of k, if A is said to be a singular matrix?
⎝ 2 1 4 ⎠

a.8
b.10
c.7
d.-6





⎛1 2 3⎞
10.let A= ⎜4 5 0⎟ then the cofactor of matrix A is the matrix
⎝2 1 4⎠,

a.

⎛ 20 −16 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−13 12 −41 ⎠

b.
⎛ 20 −16 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−13 12 −3 ⎠

c.
⎛ 20 −16 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−15 12 −3 ⎠

d.
⎛ 20 −12 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−13 12 −41 ⎠
1
Expert's answer
2019-04-04T11:46:42-0400

9. Solution: The matrix A is said to be a singular, if

"det(A)=0."

Hence 


"0=det(A)=\begin{vmatrix}
 2 & 0 & 1 \\
k & 2 & 3 \\
2 & 1 & 4 \end{vmatrix} =2\cdot 2\cdot 4 +2 \cdot 0 \cdot 3+k \cdot 1 \cdot 1 -""-2 \cdot 2 \cdot 1- k \cdot 0 \cdot 4 -2 \cdot 1 \cdot 3=k+6,"

i.e. "k=-6".

Answer: d. -6


10. Solution: First find the cofactor of each element of matrix A.


"A_{11}={(-1)}^{1+1} \begin{vmatrix}
5 & 0 \\
1 & 4 \end{vmatrix}
=20; \ A_{12}={(-1)}^{1+2} \begin{vmatrix}
4 & 0 \\
2 & 4 \end{vmatrix}=-16;""A_{13}={(-1)}^{1+3} \begin{vmatrix}
4 & 5 \\
2 & 1 \end{vmatrix}
=-6;"

"A_{21}={(-1)}^{2+1} \begin{vmatrix}
2 & 3 \\
1 & 4 \end{vmatrix}
=-5; \ 
A_{22}={(-1)}^{2+2}\begin{vmatrix}
1 & 3 \\
2 & 4 \end{vmatrix}
=-2;"


"A_{23}={(-1)}^{2+3}\begin{vmatrix}
1 & 2 \\
2 & 1 \end{vmatrix}
=3;"

"A_{31}={(-1)}^{3+1} \begin{vmatrix}
2 & 3 \\
5 & 0 \end{vmatrix}
=-15; \ 
A_{32}={(-1)}^{3+2} \begin{vmatrix}
1 & 3 \\
4 & 0 \end{vmatrix}
=12;""A_{33}={(-1)}^{3+3}\begin{vmatrix}
1 & 2 \\
4 & 5 \end{vmatrix}
=-3."

Therefore, the cofactor matrix "A^C" of the matrix "A" is equal to


"A^C=\begin{pmatrix}
20 & -16 & -6 \\
-5 & -2 & 3 \\
-15 & 12 & -3 \end{pmatrix}."

Answer: c. "\begin{pmatrix}

20 & -16 & -6 \\

-5 & -2 & 3 \\

-15 & 12 & -3 \end{pmatrix}" .